April 28, 2011 FDC meeting
- Production (Dave)
- Lessons learned from the first weeks of production
- Discoloration spots: report from Olga Trofimova
- Engineering status (Bill)
- Chamber testing
- Calibration and resolution studies (Simon) logbook page 557
- Other: talks at the collaboration meeting
Participants: Eugene, Bill, Glenn, Chris, Roger, Simon, Beni, and Lubomir.
- Problem again on Monday with the HVAC at Blue Crab: same transformer failed. It will be replaced with one with higher power, but the origin of the problem is not clear.
- Two cathodes are glued to the frames, including the flaps. Next steps: tensioning/gluing the mylar and gluing the mylar flaps on top of the copper flaps. We have to decide what should be the tension on the mylar: it's not as critical as for the kapton foil. For gluing the mylar flaps Bill suggested using conductive epoxy only in the area overlapping with the copper flaps, and regular epoxy over the g10 frame. Mark made a template needed to make holes on the mylar.
- One wire frame was strung and now they are measuring the wire positions. The new laser sensor that was mounted on the system has higher resolution but there was not enough light reflecting from the wire. We replaced it with the old sensor that, after some tuning, is being used with the old software for position measurements. The CC wants us to give them the computer used with the wire stringing system for repairing the OS. Eugene: this is critical for the production and we don't want to spend even an hour on this now.
- Lessons learned from the production so far: Lubomir will start collecting corrections to the existing procedures based on this experience. In particular we discussed the potential problem due to the extra epoxy on the inner edge of the cathode frame that can puncture the cathode. Bill suggested whipping with ethanol the epoxy from the katon right after gluing.
- A report from Olga (linked above) about the discoloration spots on the cathodes: there's chlorine but also sulfur at some places. According to Roger Allflex doesn't use ferric chloride for etching. Therefore the origin of both Cl and S is not clear. Roger will ask Allflex for samples of water that is used there. We also got foil samples from Allflex not wrapped with paper that we are going to test to exclude the paper as a possible origin. In any case the resistivity is not affected by these spots and they are not reactive, so we decided to continue with the foil production.
- Chris: got an answer from the company for the stuffing of the first 100 rigid-flexes. Bill and Lubomir: we need 18-20 for the first chamber next week; Chris will talk to them.
- Beni made pictures (pages 561-563) of the pre-amps that we use now on the prototype and then verified with Fernando the resistor settings. So we have for the cathodes 2.6mV/fC (high gain, RC=1, DR at 5% = 130fC or 338mV) and for the anodes (with and without discriminator) 0.77mV/fC (low gain, RC=0, DR at 5% 260fC or 200mV). The dynamic ranges we observe correspond to these values. Before, there was some confusion, we assumed we had 0.62mV/fC for the anode cards.
- Bill can't find a good solution for storing the cathode foils. The jig plate for the package assembling is in the machine shop.
- Simon reported his results about strip resolution and calibration. He used a new procedure (better pedestal subtraction), explained in pages 557-560 of the FDC logbook, that improved the strip resolution by ~20%. Eugene: such procedure may work in different ways on the central and side strips and one may use the sample range defined by the central strip also for the side strips. This however is difficult to implement within the fADC125 firmware.